OpenStudy (anonymous):

Could someone please help? Will give out medal! Find f(5) for this sequence: f(1) = 2 and f(2) = 3, f(n) = f(1) + f(2) + f(n - 1), for n > 2. f(5) = ______

OpenStudy (anonymous):

be specific about what kind of series are you asking?

OpenStudy (anonymous):

Thats everything the question asks

Parth (parthkohli):

Find f(3), then f(4) and then f(5).

OpenStudy (anonymous):

are there any options....? f(5)=2048

OpenStudy (anonymous):

Its a fill in the blank question

Parth (parthkohli):

Nope... f(3) = f(1) + f(2) + f(2) = 2 + 3 + 3 = 8 f(4) = f(1) + f(2) + f(3) = 2 + 3 + 8 = 13 f(5) = f(1) + f(2) + f(4) = 2 + 3 + 13 = 18

OpenStudy (anonymous):

yeah right dude.. that's the answer 18

OpenStudy (anonymous):

So you know that f(1) = 2 and f(2) =3 In order to calculate f(5) you say that n=5 so : f(5) = f(1) + f(2) + f(4) You don't know the value of f(4), and you must calculate it : f(4) = f(1) + f(2) + f(3) You dont know the value of f(3) , so you need to calculate it also: f(3) = f(1) + f(2) + f(2) = 2 + 3 + 3 =8 So f(4) = 2 + 3 + 8 = 13 AND So f(5) = 2 + 3 + 13 = 18